OpenStudy (credmond):

Solve the given equation. (Enter your answers as a comma-separated list. Let k be any integer.) 3tan^3 θ = tan θ

OpenStudy (psymon):

Well, let's do some algebra on this thing and see where it takes us: \[3\tan ^{3}\theta = \tan \theta \] Divide by 3tan(theta) \[\tan ^{2}\theta = \frac{ 1 }{ 3 }\] Now we'll square root both sides. \[\tan \theta = \pm \sqrt{\frac{ 1 }{ 3 }} \] This can be simplified to: \[\tan \theta = \pm \frac{ \sqrt{3} }{ 3 }\] Think you could take it from there?

OpenStudy (cggurumanjunath):

tan(30 deg)=1/((sqrt)3) theta =30deg.